OpenStudy (anonymous):

Help with the following ecuation. |2x+5|-|x-4|=|x^2-1| |a| means absolute value. Thanks.

hartnn (hartnn):

hi :) if you have tried it, tell me where you are stuck, we can take it from there....else it will take some time.

OpenStudy (anonymous):

well, i've tried to elevate both sides to square in order to eliminate de absolute value, but then i get horrible 8 grade ecuation. So i'm lost.

hartnn (hartnn):

oh, when i square both sides, i just get a 4th degree eq.

hartnn (hartnn):

\(|2x+5|-|x-4|=|x^2-1| \\ \sqrt{(2x+5)^2}-\sqrt{(x-4)^2}=\sqrt{(x^2-1)^2}\) because \(|a|=\sqrt{a^2}\) now square both sides...

OpenStudy (anonymous):

Of course the firts time you'll get a 4th degree eq, but you need to square again to eliminate all the abs value. \[(2x+5)^2-(x-4)^2+2*|2x+5||x-4|=(x^2-1)^2\] If you expand that, you'll get \[2*|2x+5||x-4|=x^4-7x^2-12x-40\] So i think this method so much tricky.

hartnn (hartnn):

if i square this : \(|2x+5|-|x-4|=|x^2-1| \\ \sqrt{(2x+5)^2}-\sqrt{(x-4)^2}=\sqrt{(x^2-1)^2}\) i get \((2x+5)^2+(x-4)^2-2 \sqrt{(2x+5)^2}\times \sqrt{(x-4)^2}=(x^2-1)^2 \\(2x+5)^2+(x-4)^2-2 {(2x+5)}\times {(x-4)}=(x^2-1)^2 \)

OpenStudy (anonymous):

hmmm. why \[2\sqrt{(2x+5)^2}*\sqrt{(x-4)^2}\] Becomes \[2(2x+5)(x-4)\] It shouldn't be \[2*\left| 2x+5 \right|\left| x-4 \right|\] ?

hartnn (hartnn):

yes, sorry...i am incorrect.

hartnn (hartnn):

i'll try to think of other ways...

hartnn (hartnn):

instead of squaring, lets take 2 cases everytime, \(x-4 =|2x+5|-|x^2-1| \quad or \quad x-4=|x^2-1|-|2x+5|\) do this repeatedly

hartnn (hartnn):

for first one \(|2x+5|=x-4+|x^2-1| \implies 2x+5=x-4+|x^2-1| \\ or, \: 2x+5=4-x-|x^2-1|\)

hartnn (hartnn):

getting what i am trying to do ?

OpenStudy (anonymous):

Yes, and we'll get 8 possible combinations.

hartnn (hartnn):

better than solving 8th degree equation...

hartnn (hartnn):

it would take time to solve 8 quadratic equations and check whether that particular value is actually the root of original equation or not.....but i can't find an easier way....can you continue this way? or still need help with this method ?

OpenStudy (anonymous):

This is better than squaring thinks. But happens that some values don't match with the original ecuation. Anyway, the answer it´s supposed to be x=0, and one combination provides this value. So we have to test every results and discard. I think i get it. Thanks.

hartnn (hartnn):

hmm...welcome ^_^